OpenStudy (anonymous):

sin((11π)/(12)) find the exact value

OpenStudy (anonymous):

YOUR ANSWER IS : 0.05

OpenStudy (anonymous):

thats not a choice.

OpenStudy (anonymous):

it suppose to be sqrt something and sqrt something over 4.

OpenStudy (anonymous):

WHY?

terenzreignz (terenzreignz):

Use this simple fact to your advantage~ \[\Large \frac{11\pi}{12}=\frac{2\pi}{3}+\frac{\pi}{4}\]

terenzreignz (terenzreignz):

And, of course, this identity... \[\Large \sin(\color{red}\alpha+\color{blue}\beta) = \sin(\color{red}\alpha)\cos(\color{blue}\beta) + \cos(\color{red}\alpha) \sin (\color{blue}\beta)\]

OpenStudy (anonymous):

so would it be sqrt 2 -sqrt6 4?

terenzreignz (terenzreignz):

lol I actually don't know, I was hoping you'd plug in for me :)

terenzreignz (terenzreignz):

If we work this out, we'll get to the answer.

OpenStudy (anonymous):

i dont know what to do.

OpenStudy (anonymous):

YOUR ANS IS: 0.07 GOOD LUCK.

terenzreignz (terenzreignz):

Okay, to begin with, \[\Large \sin\left(\frac{11\pi}{12}\right)= \sin\left(\color{red}{\frac{2\pi}{3}}+\color{blue}{\frac{\pi}{4}}\right)\]

terenzreignz (terenzreignz):

And use this formula \[\Large \sin(\color{red}\alpha+\color{blue}\beta) = \sin(\color{red}\alpha)\cos(\color{blue}\beta) + \cos(\color{red}\alpha) \sin (\color{blue}\beta)\]